OpenStudy (anonymous):

Solve for x. |x| – 7 = 8 A. x = –1 or 1 B. x = 1 C. x = –15 or 15 D. x = 15

OpenStudy (paki):

|x| – 7 = 8 we will get two equations here... ok x-7=8 and x-7=-8 x=8+7 and x= -8+7 x= 15 and x=-1

OpenStudy (anonymous):

thank you soo much!!!

OpenStudy (paki):

plesaure